OpenStudy (ssprague_77):

7(x - 3) = 4(x + 5) solve for x @563blackghost please help

563blackghost (563blackghost):

@AllauhAkbar that is not related to the question.

OpenStudy (ssprague_77):

What is it? My laptop won't let me open it.

563blackghost (563blackghost):

First we need to distribute the real number into the equation in the parenthesis.... \(\huge{7x-21=4x+20}\) Next we would get the real numbers to one side so we would add 21 to cancel it out and what we do to one side we do to the other... \(\huge{7x(-21+21)=4x+20+21}\) What would that simplify to?

563blackghost (563blackghost):

It was a video.

563blackghost (563blackghost):

I don't see on how it can be helpful.

OpenStudy (ssprague_77):

I did the first step, then it would come to 7x=4x+20+21 and that would be 4x=42.

OpenStudy (ssprague_77):

and i gtg schools out

563blackghost (563blackghost):

First step is correct but it would not be `4x=42`..... \(\huge{7x=4x+21+20 \rightarrow 7x=4x+41}\) In which we need to get the variable to one side so we subtract 4x from both sides... \(\LARGE{7x-4x=(4x-4x)+41} \rightarrow 3x=41\) Next we would divide by 3 to isolate the variable.... \(\huge{\frac{3x}{3}=\frac{41}{3}}\)